OpenStudy (anonymous):

how would you solve (tan^2)(1/sec^2)+1/sin in terms of sin

zepdrix (zepdrix):

Final answer in terms of sines? Hmmm, so do you remember your identities? :) sec=? tangent=? In terms of sines and cosines.

OpenStudy (anonymous):

yes sec=1/cos tan=sin/cos or \[\sin^{2}+\cos ^{2}\]

zepdrix (zepdrix):

yes we'll use that information. So let's see, what do we have when you make the conversions? :D You should have some nice cancellations.